Please give the answer according to the choices given 38-40
Co 38. Does Christine now reject or fail to reject (FTR) Ho and for what reason? FTR because isl<\td b) reject because p-value< α reject because s) reject because p-value >a c) FTR because p-value<a weighs 23 data, and determine the 39. What test do they use and why? a) t-test because its a sample s.d. c) z-test because its a large enough sample d) z-test because its a sample s.d. b) t-test because the sample is small e) none of these 40. The test is not significant at α-.01. However, using the p-value, how confident are we of the test result? a) 95.24% b) 96.88% c) 98.12% d) 99.06% e) none of these When Europeans first came to Australia, they thought that Koalas looked like bears and in fact that they were bears so they were often called Koala bears. Koalas are not bears-they are marsupials and their correct name is Koalas Though often called the koala bear, this cuddly animal is not a bear at all; it is a marsupial, or pouched mammal. After giving birth, a female koala carries her baby in her pouch for about six months wwwtoda
Answer #1

38) For question number 38 data may be given on the back side of the page. Still I conclude on the basis of the given data on this page.

Given data

n = 23

mean \bar{x} = 504.09 mg, sample s.d = 40 mg

level of significance = 0.01

test statistic t = \frac{\bar{x}-\mu }{s/\sqrt{n}}

in this problem \mu value is not provided. take that value and find the test statistic.

and calculate critical value tc if calculated test statistic |t| > |tc| then reject null hypothesis or else if p-value is less than \alpha then also reject null hypothesis.

39) We can use t-test because sample standard deviation is known and population standard deviation is unknown(may be given in data for question number 38, if it is given then your option will be b). And also the sample size is small.